# AlloyASG: Alloy Predicate Code Representation as a Compact Structurally Balanced Graph

### v3

@@ Abstract.
In the program analysis and automated bug-fixing fields, it is common to create an abstract interpretation of a program’s source code as an Abstract Syntax Tree (AST), which enables programs written in a high-level language to have various static and dynamic analyses applied. However, ASTs suffer from exponential growth in their data size due to the limitation that ASTs will often have identical nodes separately listed in the tree. To address this issue, we introduce a novel code representation schema, Complex Structurally Balanced Abstract Semantic Graph (CSBASG), which represents code as a complex-weighted directed graph that lists a semantic element as a node in the graph and ensures its structural balance for almost finitely enumerable code segments, such as the modeling language Alloy. Our experiment ensures that CSBASG provides a one-on-one correspondence of Alloy predicates to complex-weighted graphs. We evaluate the effectiveness and efficiency of our CSBASG representation for Alloy models and identify future applications of CSBASG for Alloy code generation and automated repair.
Alloy, Abstract Semantic Graphs, Program Representation, Automatic fixing, Structural balance, Complex-weighted graphs
@@ 1. Introduction
Alloy (alloy, 16) is a declarative, relational logic-based modeling language for describing structural and dynamic properties of a system’s design. Since Alloy is used to verify software system designs (Zave15, 40, 5, 34, 8), and to perform various forms of analyses over the corresponding implementation, including deep static checking (JacksonVaziri00Bugs, 17, 10), systematic testing (MarinovKhurshid01TestEra, 22), data structure repair (ZaeemKhurshidECOOP2010, 39), automated debugging (GopinathETALTACAS2011, 12) and to synthesize security attacks (websecurity, 2, 23, 29), the consequences could be catastrophic if an Alloy model is faulty yet successfully compiles. Several iterative search methods based on ASTs have already been implemented for automated repair and partial generation of Alloy models (amra, 32, 14, 42, 7). However, any iterative search methods for mutants or AST elemental changes that fix the Alloy code segments require expansive storage and have intractable search spaces due to the exponential growth of the code segment length, limiting their computational efficiencies.
@@
In recent years, there have been several bodies of work that focus on automatically repairing programming code using machine learning techniques (mlfix1, 30, 20, 13, 3, 38, 15, 28). Such works provide valuable inspiration to us since we have a dataset of real-world fixing pairs consisting of faulty predicates written by the students and their correct counterparts (alloy4fun, 21), effectively leading to a set of pairs of inputs and ground truths. Nonetheless, almost all the sequence-based algorithms treat the code segments as natural language sequences with a black-box function trained for faulty code as input and corrected code as output. These approaches ignore the strict logic of the programming languages, severely hindering the explainability of the process.
@@
To create a machine learning algorithm that is aware of the logic of the programming language we are generating code in, we need a logical structure-aware, vectorized, machine-and-human-readable, and computationally efficient code representation that improves upon the AST adjacency matrix representation that grows exponentially in size. Some of the representation learning methods (code2seq, 4, 25) train functions that output vectors from the AST, yet the nature of their dependency on the control flow paths in the AST of those general-purpose languages renders their methodologies implausible for a specification language such as Alloy, which is declarative and lacks control flow. Other methods (astnn, 41, 33, 35) try to focus on the raw AST for a universal representation, yet those algorithms still suffer from rapid growth in the size of the AST. Therefore, this paper aims to create a structure that generates a compact representation of the code segment, yet still gives the same complete information as the raw AST. Recent work created an Abstract Semantic Graph (ASG) (DuffyASG, 9), which is close to meeting our data representation needs, but the original designation lacks a numerical representation that provides an adjacency or Laplacian matrix that could be used in popular graph machine learning algorithms (gcn1, 24, 11, 18, 31, 27). Therefore, in this paper, we introduce a matrix representation of the ASG with the following objectives:
@@
    better performance in compactness and possible space-saving than the raw AST representation;
@@
    allowing a direct comparison of the difference between the pair of code segments before and after manual fixing.
@@
One of the advantages of the graph is the uniqueness of the semantics. For two semantically different terms in an AST, e.g., a variable or a unary or binary operator, there will only be one corresponding vertex in the CSBASG. This intuitively represents the code elements and their relationships with a reduced space cost. Besides, for a fixing pair of code segments consisting of the faulty code and its correction, CSBASG could align the modifications by the common node signatures shared by the pair and compare their structural differences.
@@
Open Source: We release our source code to produce https://github.com/AlexandervonWu/AlloyASG-Release.
@@ 2. Background
In this section, we describe key components of Abstract Semantic Graphs and the Alloy modeling language.
@@ 2.1. Abstract Syntax Tree and Abstract Semantic Graph
We use the common notation of the abstract syntax tree (AST) for the raw representation of the syntactic structure of the Alloy source code (ast, 6). We follow the formal definitions from (autoinference, 19):
@@
(autoinference, 19) A context-free grammar (CFG) $G\equiv(N,\Sigma,R,s)$ is a tuple where $N$ is the set of nonterminals, $\Sigma$ is the set of terminals, $R$ is the set of production rules $a\to b_{1}b_{2}...b_{N}$ where $a\in N$ is a nonterminal and $b_{i}\in N\cup\Sigma$, and $s\in N$ is the starting nonterminal of the grammar. The language of $G$ is the set of strings from the starting non-terminal.
@@
(autoinference, 19) Given a context-free grammar (CFG) $G=(N,\Sigma,R,s)$, an abstract syntax tree (AST) $T\equiv(G,X,r,\xi,\sigma)$ is a tree with a set of nodes $X$, a root node $r\in X$, a mapping $\xi:X\to\mathcal{P}(X)$ from each of the nodes to a subset of nodes as its children and each child set $\forall x\in X:\xi(x)$ is either empty or well-ordered, and $\sigma:X\to(N\cup\Sigma)$ maps each node to its corresponding label as a terminal or nonterminal word in the CFG.
According to (DuffyASG, 9), an abstract semantic graph (ASG) resolves the branches defining the properties of a node in the AST to its original definitions. But here, to create a more compact form of ASG specifically for Alloy, we use a more radical definition that combines every syntactically and semantically equivalent node as below:
@@
Note that since $\mathcal{E}$ is an arbitrary set of edges in forms of $(v_{i},v_{j},w_{ij})$, an AST or empirical construction of ASG could have more than one set of $\{w_{ij}\}$ to represent. If the weight function $w:\mathcal{V}\times\mathcal{V}\to\mathbb{C}$ ensures that $w_{ij}\neq 0$ if and only if there exists an edge from $v_{i}$ to $v_{j}$, and $w_{ij}\neq w_{ik}$ for any $j\neq k$ be children of $i$, either at the same order of execution (children under the same node in the corresponding AST) or at the different order (if a nonterminal presents more than once in the AST and each time comes with a set of children). There could also be multiple AST links connecting two ASG nodes, and we will discuss how we form a weight value for this case in Section 3.
@@ 2.2. Complex-weighted Graph and its Structural Balance
Since Definition 3 gives a complex-weighted graph and our requirement for a vectorization of such ASG, the construction of the graph requires the edge values specifying the relationships between any of the two nodes. So we borrow the structural balance concept from (sbg, 36):
@@
(sbg, 36) For a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, let $a_{ij}=w_{ji}$ be the entries of the adjacency matrix $A$ where $|a_{ij}|>0$ indicates there exists an edge between $v_{j}$ and $v_{i}$, then the graph $\mathcal{G}$ is said to be structurally balanced if all the entries of its adjacency matrix $A=[a_{ij}]^{N\times N}$ satisfies $a_{ij}\equiv|a_{ij}|\angle\theta_{ij}=|a_{ij}|\angle(\theta_{i}-\theta_{j})$, where $\theta_{1},...,\theta_{N}\in(-\pi,\pi)$ is called the signatures of the nodes $v_{1}...v_{N}$, respectively.
In the ASG context, intuitively, the signatures of the nodes $\theta_{1},...,\theta_{n}$ could encode the syntactic and semantic properties of each terminal or nonterminal node and the magnitudes of each edge $|a_{ij}|_{i,j\in\{1...N\}}$ encodes their relative properties, such as if a node is the left (1st) or right (2nd) child of a binary operator. We chose this concept for its mathematical solid properties. Let $d_{i}=\sum_{j\in 1}^{N}|a_{ij}|$ which is the in-degree of node $v_{i}$ and let $D=diag(d_{1},...,d_{N})$ be the diagonal matrix giving the in-degrees of each node on their corresponding rows, and define $L=D-A$ as the Laplacian matrix, then we have the lemma below (sbg, 36):
Lemma 0 ().
(sbg, 36, 37)
The following are equivalent:
(1) Complex weighted graph $\mathcal{G}(A)$ is structurally balanced.
(2) Zero is a eigenvalue of $L$ with eigenvector $\bm{\zeta}=[1\angle\theta_{1},\dots,1\angle\theta_{N}]^{T}$.
(3) $D_{\zeta}:=diag(\bm{\zeta})$ such that $\widehat{A}=D_{\zeta}^{-1}AD_{\zeta}$ is nonnegative and $\widehat{A}=[|a_{ij}|]_{N\times N}$.
(4) $D_{\zeta}:=diag(\bm{\zeta})$ such that $\widehat{L}=D_{\zeta}^{-1}LD_{\zeta}$ has a zero eigenvalue with an eigenvector being $\bm{1}$, where $\widehat{L}=D-\widehat{A}$.
Proof.
See proof of Lemma 1 in (sbgcomplete, 37). ∎
Notably, in the notation of (sbg, 36, 37), each column specifies the outer edges of a node, and the row-based adjacency matrix must be transposed before acquiring the Laplacian.
@@ 2.3. Alloy
Alloy users write models that describe the properties of the system of interest. Then, the Analyzer helps the user understand their system by displaying the consequences of their properties, helping identify any missing or incorrect properties, and exploring the impact of modifications to those properties. To achieve this, the Analyzer uses off-the-shelf SAT solvers to search for scenarios, which are assignments to the sets and relations of the model such that all executed formulas hold. These scenarios are presented visually and introduce a separation of concern between writing and exploring a specification. If no such scenario can be found, the Analyzer reports that the formulas are unsatisfiable.
To highlight how modeling in Alloy works, Figure 1 depicts the base model of a classroom management system. This model is from the Alloy4Fun dataset, which is a collection of submissions made by novice users learning Alloy (alloy4fun, 21). Signature paragraphs introduce named sets and can define relations, which outline relationships between elements of sets. Line 1 introduces a named set Person and establishes that each Person atom connects to any number of (set) Person atoms through the Tutor relation and each Person atom connects to any number of (set) Class atoms through the Teaches relation. Line 5 introduces the named set Group, which contains no relations. Line 6 introduces the named set Class and states that each class has a set of people assigned to a group using the ternary relational Groups. Lines 7 and 8 introduce the named sets Teacher and Student as subsets (in) of Person.
Predicate paragraphs introduce named first-order, linear temporal logic formulas that can be invoked elsewhere. Figure 1 (b) depicts the oracle for exercise inv15. The predicate inv15oracle uses universal quantification (‘all’), set multiplicity (‘some’), set intersection (‘&’), transitive closure (‘^’), and relational join (‘.’) to express that for every person, there is an intersection between the set Teacher and the set of Tutors reachable by one or more invocation of Tutor’s relation. The student submission in Figure 1 (c) is faulty, and and is overconstrained, which means the faulty formula prevents instances that should have been generated, and underconstrained, which means the faulty formula allows both instances that should have been prevented. The student attempted to use nested quantification to explicitly outline that a teacher should be reachable in 1 to 3 traversals down the Tutor’s relation.
2.4. Alloy Grammar
@@
We consider the Alloy grammar (alloy, 16) before constructing the graph representation. The basic blocks under the overall root of an Alloy model, ModelUnit, consists of a set of declarations, such as signatures as basic types, functions as processes, and predicates as logical arguments. Our dataset Alloy4Fun (alloy4fun, 21) covers only simple predicates without references to external libraries; therefore, creating representations for those code segments included in the dataset is an ideal starting point. Figure 2 gives an example of an Alloy predicate in the dataset that contains 32 AST nodes. From this representation, it is obvious that the raw AST representation has been relatively huge and complicated, even with a short segment of code enveloped in a single predicate, mostly due to the redundancy of the semantically identical nodes presenting in the AST for their multiple appearances in the original code. In the given example, there are three binary operators ‘‘in’’, and 6 dots (JOIN) as operators; these would all be treated as distinct nodes in a raw AST representation.
A predicate defines a scope that contains an expression while itself is a solid, invariant root for any fixing pairs. We can easily enumerate all possible nodes under an expression (or formula): the finitely enumerable types of expressions or formulae, plus variable declarations that only occur under the quantifiers of first-order logic. In this paper, we assign them fixed signatures, yet an adaptive signature system could be used in case of an extension of the CSBASG into either more types of nodes or extra optimizations.
Another property of Alloy is that almost all expression or formula nodes have a finite and fixed number of children, for instance, 1 for unary operators, 2 for binary operators, and 3 for the if-then-else sentences. In these cases, there could be a trivial, polynomial-based implementation of the $|w_{ij}|$ values for the edges between such a node $v_{i}$ and one of its children $v_{j}$. Even so, there are some exceptions, such as a function call (having an indefinite number of parameters), a list expression or formula (for example, the consecutive logical ANDs and ORs for a conjunctive or disjunctive normal form), or a set of variable declarations for the same quantifier. We will handle them on a case-by-case basis in Section III.

### v4

@@ Abstract.
Writing declarative models has numerous benefits, ranging from automated reasoning and correction of design-level properties before systems are built to automated testing and debugging of their implementations after they are built. Unfortunately, the model itself needs to be correct to gain these benefits. Alloy is a commonly used modeling language that has several existing efforts to repair faulty models automatically. Currently, these efforts are search-based methods that use an Abstract Syntax Tree (AST) representation of the model and do not scale. One issue is that ASTs themselves suffer from exponential growth in their data size due to the limitation that ASTs will often have identical nodes separately listed in the tree. To address this issue, we introduce a novel code representation schema, Complex Structurally Balanced Abstract Semantic Graph (CSBASG), which represents code as a complex-weighted directed graph that lists a semantic element as a node in the graph and ensures its structural balance for almost finitely enumerable code segments. We evaluate the efficiency of our CSBASG representation for Alloy models in terms of it’s compactness compared to ASTs, and we explore if a CSBASG can ease the process of comparing two Alloy predicates. Moreover, with this representation in place, we identify several future applications of CSBASG, including Alloy code generation and automated repair.
Alloy, Abstract Semantic Graph, Program Representation, Automatic Fixing, Structural Balance, Complex-weighted Graph
@@ 1. Introduction
Alloy (Jackson, 2002) is a declarative, relational logic-based modeling language for describing the structural and dynamic properties of a system. Since Alloy is used to verify software system designs (Zave, 2015; Bagheri et al., 2018; Wickerson et al., 2017; Chong et al., 2018), and to perform various forms of analyses over the corresponding implementation, including deep static checking (Jackson and Vaziri, 2000; Galeotti et al., 2013), systematic testing (Marinov and Khurshid, 2001), data structure repair (Zaeem and Khurshid, 2010), automated debugging (Gopinath et al., 2011) and to synthesize security attacks (Akhawe et al., 2010; Nelson et al., 2010; Trippel et al., 2019), the consequences could be catastrophic if an Alloy model is faulty yet successfully compiles. Several iterative search methods based on ASTs have already been implemented for automated repair and partial generation of Alloy models (Wang et al., 2018a; Gutiérrez Brida et al., 2021; Zheng et al., 2022; Cerqueira et al., 2022). However, any iterative search methods for mutants or AST elemental changes that fix the Alloy code segments require expansive storage and have intractable search spaces due to the exponential growth of the code segment length, limiting their computational efficiencies.
@@
In recent years, there have been several bodies of work that focus on automatically repairing code using machine learning techniques (Tufano et al., 2019; Long and Rinard, 2016; Gupta et al., 2017; Alhefdhi et al., 2020; Yang et al., 2020; Huang et al., 2021; Sobania et al., 2023). Such works provide valuable inspiration to us since we have a dataset of real-world fixing pairs consisting of faulty predicates written by students and their correct counterparts (Macedo et al., 2019), effectively leading to a set of pairs of inputs and ground truths. Nonetheless, almost all the sequence-based algorithms treat the code segments as natural language sequences with a black-box function trained for faulty code as input and corrected code as output. These approaches ignore the strict logic of the target programming language, severely hindering the explainability of the process and limiting their applicability to modeling languages like Alloy.
@@
To create a machine learning algorithm that is aware of the logic of the programming language we are generating code in, we need a logical structure-aware, vectorized, machine-and-human-readable, and computationally efficient code representation that improves upon the AST adjacency matrix representation that grows exponentially in size. Some of the representation learning methods (Alon et al., 2019; Paassen et al., 2021) train functions that output vectors from the AST, yet the nature of their dependency on the control flow paths in the AST of those general-purpose languages renders their methodologies implausible for a specification language such as Alloy, which is declarative and lacks control flow. Other methods (Zhang et al., 2019; Wang et al., 2022; Wu et al., 2022) try to focus on the raw AST for a universal representation, yet those algorithms still suffer from rapid growth in the size of the AST. Therefore, this paper aims to create a structure that generates a compact representation of the code segment yet still gives the same complete information as the raw AST.
Recent work created an Abstract Semantic Graph (ASG) (Duffy and Malloy, 2012), which is close to meeting our data representation needs. Still, the original designation lacks a numerical representation that provides an adjacency or Laplacian matrix that could be used in popular graph machine learning algorithms (Niepert et al., 2016; Gao et al., 2018; Li et al., 2018; Wang et al., 2017; Schlichtkrull et al., 2017). Therefore, in this paper, we introduce a matrix representation of the ASG with the following objectives:
@@
    Better performance in compactness and the potential to save space compared to the raw AST representation;
@@
    Allow a direct comparison of the differences between code segments before and after applying a fix.
@@
One of the advantages of this graph representation is the uniqueness of the semantics. For two semantically different terms in an AST, e.g., a variable or a unary or binary operator, there will only be one corresponding vertex in the CSBASG. This intuitively represents the code elements and their relationships with a reduced space cost. Besides, for a fixing pair of code segments consisting of the faulty code and its correction, the CSBASG could align the modifications by the common node signatures shared by the pair and compare their structural differences.
@@
Open Source: We release our source code at https://anonymous.4open.science/r/AlloyASG-Release-CF0F/ (to be replaced with public GitHub in the final version).
Future Directions: We present a series of research directions now enabled due to the representation of Alloy models with a CSBASG.
@@ 2. Background
This section describes key components of Abstract Semantic Graphs and the Alloy modeling language.
@@ 2.1. Abstract Syntax Tree and Abstract Semantic Graph
We use the common notation of the abstract syntax tree (AST) for the raw representation of the syntactic structure of the Alloy source code (Baxter et al., 1998). We follow the formal definitions from (Long et al., 2017):
@@
(Long et al., 2017) A context-free grammar (CFG) $G\equiv(N,\Sigma,R,s)$ is a tuple where $N$ is the set of nonterminals, $\Sigma$ is the set of terminals, $R$ is the set of production rules $a\to b_{1}b_{2}...b_{N}$ where $a\in N$ is a nonterminal and $b_{i}\in N\cup\Sigma$, and $s\in N$ is the starting nonterminal of the grammar. The language of $G$ is the set of strings from the starting non-terminal.
@@
(Long et al., 2017) Given a context-free grammar (CFG) $G=(N,\Sigma,R,s)$, an abstract syntax tree (AST) $T\equiv(G,X,r,\xi,\sigma)$ is a tree with a set of nodes $X$, a root node $r\in X$, a mapping $\xi:X\to\mathcal{P}(X)$ from each of the nodes to a subset of nodes as its children and each child set $\forall x\in X:\xi(x)$ is either empty or well-ordered, and $\sigma:X\to(N\cup\Sigma)$ maps each node to its corresponding label as a terminal or nonterminal word in the CFG.
According to (Duffy and Malloy, 2012), an abstract semantic graph (ASG) resolves the branches defining the properties of a node in the AST to its original definitions. But here, to create a more compact form of an ASG specifically for Alloy, we use a more radical definition that combines every syntactically and semantically equivalent node as below:
@@
Note that since $\mathcal{E}$ is an arbitrary set of edges in forms of $(v_{i},v_{j},w_{ij})$, an AST or empirical construction of ASG could have more than one set of $\{w_{ij}\}$ to represent. If the weight function $w:\mathcal{V}\times\mathcal{V}\to\mathbb{C}$ ensures that $w_{ij}\neq 0$ if and only if there exists an edge from $v_{i}$ to $v_{j}$, and $w_{ij}\neq w_{ik}$ for any $j\neq k$ be children of $i$, either at the same order of execution (children under the same node in the corresponding AST) or at the different order (if a nonterminal presents more than once in the AST and each time comes with a set of children). There could also be multiple AST links connecting two ASG nodes, and we will discuss how we form a weight value for this in Section 3.
@@ 2.2. Complex-weighted Graph and its Structural Balance
Since Definition 3 gives a complex-weighted graph and our requirement for a vectorization of such ASG, the construction of the graph requires the edge values specifying the relationships between any of the two nodes. So we borrow the following structural balance concept from (Wu et al., 2023a):
@@
(Wu et al., 2023a) For a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, let $a_{ij}=w_{ji}$ be the entries of the adjacency matrix $A$ where $|a_{ij}|>0$ indicates there exists an edge between $v_{j}$ and $v_{i}$, then the graph $\mathcal{G}$ is said to be structurally balanced if all the entries of its adjacency matrix $A=[a_{ij}]^{N\times N}$ satisfies $a_{ij}\equiv|a_{ij}|\angle\theta_{ij}=|a_{ij}|\angle(\theta_{i}-\theta_{j})$, where $\theta_{1},...,\theta_{N}\in(-\pi,\pi)$ is called the signatures of the nodes $v_{1}...v_{N}$, respectively.
In the ASG context, intuitively, the signatures of the nodes $\theta_{1},...,\theta_{n}$ could encode the syntactic and semantic properties of each terminal or nonterminal node and the magnitudes of each edge $|a_{ij}|_{i,j\in\{1...N\}}$ encodes their relative properties, such as if a node is the left (1st) or right (2nd) child of a binary operator. We chose this concept for its mathematical solid properties. Let $d_{i}=\sum_{j\in 1}^{N}|a_{ij}|$ which is the in-degree of node $v_{i}$ and let $D=diag(d_{1},...,d_{N})$ be the diagonal matrix giving the in-degrees of each node on their corresponding rows, and define $L=D-A$ as the Laplacian matrix, then we have the lemma below (Wu et al., 2023a):
Lemma 0.
(Wu et al., 2023a, b)
@@
The following are equivalent:
(1):
Complex weighted graph $\mathcal{G}(A)$ is structurally balanced.
(2):
Zero is a eigenvalue of $L$ with eigenvector $\bm{\zeta}=[1\angle\theta_{1},\dots,1\angle\theta_{N}]^{T}$.
(3):
$D_{\zeta}:=diag(\bm{\zeta})$ such that $\widehat{A}=D_{\zeta}^{-1}AD_{\zeta}$ is nonnegative and $\widehat{A}=[|a_{ij}|]_{N\times N}$.
(4):
$D_{\zeta}:=diag(\bm{\zeta})$ such that $\widehat{L}=D_{\zeta}^{-1}LD_{\zeta}$ has a zero eigenvalue with an eigenvector being $\bm{1}$, where $\widehat{L}=D-\widehat{A}$.
Proof.
See proof of Lemma 1 in (Wu et al., 2023b). ∎
Notably, in the notation of (Wu et al., 2023a, b), each column specifies the outer edges of a node, and the row-based adjacency matrix must be transposed before acquiring the Laplacian.
@@ 2.3. Alloy
Alloy users write models that describe the properties of the system of interest. Then, the Analyzer helps the user understand their system by displaying the consequences of their properties, helping identify any missing or incorrect properties, and exploring the impact of modifications to those properties. To achieve this, the Analyzer uses off-the-shelf SAT solvers to search for scenarios, which are assignments to the sets and relations of the model such that all executed formulas hold. If no such scenario can be found, the Analyzer reports that the formulas are unsatisfiable.
To highlight how modeling in Alloy works, Figure 1 depicts the base model of a classroom management system. This model is from the Alloy4Fun dataset, which is a collection of submissions made by novice users learning Alloy (Macedo et al., 2019). Signature paragraphs introduce named sets and can define relations, which outline relationships between elements of sets. Line 1 introduces a named set Person and establishes that each Person atom connects to any number of (set) Person atoms through the Tutor relation and each Person atom connects to any number of (set) Class atoms through the Teaches relation. Line 5 introduces the named set Group, which contains no relations. Line 6 introduces the named set Class and states that each class has a set of people assigned to a group using the ternary relational Groups. Lines 7 and 8 introduce the named sets Teacher and Student as subsets (in) of Person.
[…]
Figure 2. Counterexample highlighting difference between inv15 and inv15oracle
Predicate paragraphs introduce named first-order, linear temporal logic formulas that can be invoked elsewhere. Figure 1 (b) depicts the oracle for exercise inv15. The predicate inv15oracle uses universal quantification (‘all’), set multiplicity (‘some’), set intersection (‘&’), transitive closure (‘^’), and relational join (‘.’) to express that for every person, there is some intersection between the set Teacher and the set of Tutors reachable from person p. Figure 1 (c) displays a faulty student submission for this exercise. The student attempts to use nested quantification to explicitly outline that a teacher should be reachable in 1 to 3 traversals down the Tutor’s relation. The student submission assumes incorrectly that a person’s tutor relation captures the set of all people who tutor $p$, but the tutor relation captures who $p$ themselves tutors. Therefore, the student submission can be corrected by inserting the transpose operator, e.g., “p.~Tutors.”
Figure 2 displays one of the counterexamples produced when the student submission is checked against the backend oracle in Figure 1 (b). The depicted scenario is valid for the oracle solution but not for student submission. To find this counterexample, an Alloy command is run that invokes the backend SAT solver. All commands require a scope, which places an upper bound of the universe of discourse. The default scope is 3, which means Alloy tries to find a counterexample highlighting that the two formulas are not equivalent using up to 3 Person, 3 Group, and 3 Class atoms.
2.4. The Alloy Grammar
@@
We first highlight aspects of the Alloy grammar (Jackson, 2002) that influence how we construct the graph representation. An Alloy model consists of a set of declarations with signatures as basic types, functions as processes, and predicates as logical arguments. Our dataset Alloy4Fun (Macedo et al., 2019) covers only simple predicates without references to external libraries; therefore, creating representations for those code segments included in the dataset is an ideal starting point. Figure 3 gives an example of an Alloy predicate in the dataset that contains 32 AST nodes. From this representation, it is obvious that the raw AST representation is relatively huge and complicated, even with a short segment of code enveloped in a single predicate, mostly due to the redundancy of semantically identical nodes that appear multiple times in the AST since they appear multiple times in the original code. In the given example, there are three subset operators “in” and 6 relational join “.” operators that are all treated as distinct nodes in the raw AST representation.
A predicate defines a scope that contains an expression while itself is a solid, invariant root for any fixing pair. We can easily enumerate all possible nodes under an expression (or formula): the finitely enumerable types of expressions or formulae, plus variable declarations that only occur under the quantifiers of first-order logic. In this paper, we assign them fixed signatures, yet an adaptive signature system could be used in the case of an extension of the CSBASG into either more types of nodes or extra optimizations.
Another property of Alloy is that almost all expression or formula nodes have a finite and fixed number of children, for instance, 1 for unary operators, 2 for binary operators, and 3 for the if-then-else sentences. In these cases, there could be a trivial, polynomial-based implementation of the $|w_{ij}|$ values for the edges between such a node $v_{i}$ and one of its children $v_{j}$. Even so, there are some exceptions, such as a function call (having an indefinite number of parameters), a list expression or formula (for example, the consecutive logical ANDs and ORs for a conjunctive or disjunctive normal form), or a set of variable declarations for the same quantifier. We will handle them on a case-by-case basis in Section 3.